How much momentum is acquired by a 75-kg sky-diver
Answer #1

step:1

Given data

mass m=75 kg

time t=2 min

step:2

use the kinematices to find the velocity driver

=>v-u=gt

=>v=9.8*2*60=1176 m/s

the velocity of driver falls v=1176 m/s

step:3

now we find the momentum

momentum P=mv=75*1176=88200 kg.m/s
Answer #2

SOLUTION :


Velocity of the driver 2 minute after jumping, v :

= g * (2 *60)

= 9.8 * 120

= 1176 m/s 


So, momentum acquired by the driver at the instant

= m * v

= 75 * 1176

= 88200 kg - m / s.  (ANSWER).

answered by: Tulsiram Garg